three spheres each of mass r/3 and radius m/2, are kept such that each touches other two. the magnitude of gravitational force is given by

Grade:9

1 Answers

PRATYUSH Khare
4 years ago
Three spheres are kept in such a way that they form an eqilateral triangle.
Using Newton's Law of gravitation,
F= GmM/r2 
F=G(R/3)2/(m/2+m/2)2........(Radius gets added)
F=GR2/9../M2
Using Resultant formula
√[F2 + F2 + 2ABcos●]
√[F2 + F2 + 2F2cos60]
= √3F2
=√3GR2/9.../M2
Can be simplified further......